# Geometry and Optimization of Shallow Polynomial Networks

### v1

@@ Abstract
We study shallow neural networks with polynomial activations. The function space for these models can be identified with a set of symmetric tensors with bounded rank. We describe general features of these networks, focusing on the relationship between width and optimization. We then consider teacher-student problems, that can be viewed as a problem of low-rank tensor approximation with respect to a non-standard inner product that is induced by the data distribution. In this setting, we introduce a teacher-metric discriminant which encodes the qualitative behavior of the optimization as a function of the training data distribution. Finally, we focus on networks with quadratic activations, presenting an in-depth analysis of the optimization landscape. In particular, we present a variation of the Eckart-Young Theorem characterizing all critical points and their Hessian signatures for teacher-student problems with quadratic networks and Gaussian training data.
@@ 1 Introduction
One of the main challenges in the theory of deep learning is to explain why neural networks can be successfully trained in practice despite the fact that their loss is highly non-convex. In recent years, researchers have argued that this behavior is mainly a consequence of overparameterization, that is, the fact that models have a large number of parameters compared to the number training samples. However, while the benefit of overparameterization has been demonstrated in many settings, precise theoretical results often rely on unrealistic assumptions, such as infinite neurons [13, 26, 36], or only apply to very restricted models, such as linear networks [27]. Additionally, several works provide sufficient conditions for the number of parameters needed for favorable optimization [18, 42], but they do not consider important information such as the training data distribution or the initialization of gradient descent.
@@
In this paper, we seek a more precise understanding of the conditions under which local minima pose a challenge to gradient descent. Our general setting is based on shallow polynomial networks, that are functions $f_{\mathcal{W}}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ of the form
$f_{\mathcal{W}}(x)=\sum_{i=1}^{r}\alpha_{i}(w_{i}\cdot x)^{d},\qquad\mathcal{W}=(\alpha,w_{1},\ldots,w_{k})\in\mathbb{R}^{r}\times(\mathbb{R}^{n})^{k},$ (1)
@@
where $d\in\mathbb{N}$. We believe that shallow polynomial networks capture many essential features of the optimization properties for more general networks — e.g., with ReLU activations — while also being very mathematically appealing. In particular, shallow polynomial networks are intimately related to symmetric tensor decompositions: the set $\mathcal{F}_{r}$ of all shallow networks of width at most $r$ can be identified with the set of symmetric tensors of rank at most $r$. Though this connection has been noted in prior works [42, 30], its full implications for optimization have not yet been studied in detail. Our focus on algebraic networks aligns with recent work investigating neural network architectures using algebraic-geometric tools [33, 32, 25, 34, 20]. It is worth noting that the fact that polynomial activations are not universal approximators does not represent a limitation, as our study is centered on non-convex optimization. Moreover, polynomial networks with varying degrees are, in fact, universal approximators.
@@
We begin by revisiting the connection between tensors and shallow networks, describing properties of tensors that offer insight into these models. A key aspect that emerges from this relationship is the existence of three distinct regimes for the network width $r$: 1) for small $r$, the set $\mathcal{F}_{r}$ of shallow networks is a low-dimensional subset of its ambient space; 2) for intermediate $r$, the set $\mathcal{F}_{r}$ is a full-dimensional (semi-algebraic) set and yet strictly contained in its ambient space; 3) for large $r$, the set $\mathcal{F}_{r}$ “fills” the space of all symmetric tensors completely. These settings lead to very different behaviors when optimizing training losses, with the filling regime being the “easiest” in a precise sense. The Alexander-Hirschovitz Theorem [2] from the theory of tensors effectively classifies the sharp transitions between these regimes for all input dimensions $n$ and activation degrees $d$.
@@
After discussing general properties of polynomial networks, we focus on teacher-student training problems. For any $r,s\geq 1$, we consider a loss function of the form
@@
[…]
@@
In the presence of symmetry, critical points of (2) were identified as examples of symmetry breaking [6] (see generally [5]) under the class of inner products considered in this work, enabling the construction of infinite families of critical points and the characterization of their Hessian spectrum [7]. On the other hand, when symmetry is absent, the optimization landscape for low-rank tensor approximation is not fully understood. Existing results only consider the Frobenius norm and are typically restricted to rank-$1$ approximation [22, 17, 4, 29]. Even in the simpler setting of low-rank matrix approximation, the behavior under weighted norms is not yet fully understood [23]. For this reason, after some general results for arbitrary activation degrees, we conduct a detailed study for the particular case of quadratic activations. Shallow networks with quadratic activations have been the subject of several papers [18, 42, 40, 21, 39]. Compared to these works, our analysis provides a more general and detailed picture of the loss landscape. In particular, we provide a characterization of all critical points and their signatures in the case of Gaussian (or rotationally invariant) data distributions. Perhaps surprisingly, we also find that the teacher-student loss for “generic” data distributions behaves very differently, and leads to many more critical points than the Gaussian case. This result illustrates potential dangers of theoretical works focusing exclusively on Gaussian data. Throughout the paper, tools of applied algebraic geometry allow us to achieve novel analysis for shallow networks.
Organization.
In Section 2, we present general properties of shallow polynomial networks. We introduce the functional space of these models and the three basic regimes (low-dimensional, thick, filling) for the network width (Section 2.1). We also discuss the parameterization of these spaces and characterize the branch locus for thick architectures with activation degree $2$ or input dimension $2$ (Proposition 6 and Proposition 7). We then consider the optimization landscape of these networks as a function of width, refining several results from [42](Proposition 10, Theorem 11Theorem 12). Moreover, while prior works argued that very wide networks have favorable landscapes, we show that bad local minima can exist for networks of any width (Example 15). Additionally, these minima may have a basin of attraction with a positive Lebesgue measure (Corollary 17). This apparent contradiction arises from differing notions of unfavorable landscapes (“bad minima” and “spurious valleys”). In Section 3, we introduce teacher-student problems, and relate them to the problem of low-rank tensor approximation with respect to non-standard norms. We prove that a data distribution induces a quadratic form whose entries are given by moments of the distribution, and provide examples of norms arising from different classes of distributions (Section 3.1). We revisit the general problem of finding the nearest point on a manifold or variety, adapting classical results on focal locus [37] (or ED discriminants [16]) to the case of varying metrics/data distributions (Theorem 25, Theorem 26). We also present examples of discriminants with computations performed using Macalauy2 (Section 3.3). In Section 4 we present a detailed case-study for networks with quadratic activations. We specialize some results from the previous sections by describing the relation between critical points in function space and parameter space (Section 4.1), and giving simpler expressions for induced inner products (Section 4.2). We then present a version of the Eckart-Young Theorem which describes all critical points and their indices for teacher-student problems with quadratic networks and Frobenius (Theorem 34) or Gaussian (Theorem 35) norms. Finally, we prove that for other norms, including those arising from general i.i.d. distributions, there are generally exponentially more critical points (Theorem 36).
@@
The Frobenius inner product of two tensors $T$ and $S$ of the same shape is given by $\langle T,S\rangle_{F}=\sum_{(i_{1},\ldots,i_{d})}T_{i_{1}\ldots i_{d}}S_{i_{1}\ldots i_{d}}$. If $T=v_{1}\otimes\cdots\otimes v_{d}$ and $S=w_{1}\otimes\cdots\otimes w_{d}$, then $\langle T,S\rangle_{F}=(v_{1}\cdot w_{1})\cdots(v_{d}\cdot w_{d})$, and in particular, $\langle v^{\otimes d},w^{\otimes d}\rangle_{F}=(v\cdot w)^{d}$. A tensor $T\in\mathbb{R}^{n}\otimes\cdots\otimes\mathbb{R}^{n}$ (with $d$ factors) is symmetric if it is invariant under any permutation of indices, i.e., $T_{i_{1},\ldots,i_{d}}=T_{i_{\sigma(1)},\ldots,i_{\sigma(d)}}$ for any $\sigma\in\mathfrak{S}_{d}$, the symmetric group of degree $d$. We denote the space of symmetric tensors of order $d$ on $\mathbb{R}^{n}$ by $\text{Sym}^{d}(\mathbb{R}^{n})$, a vector space of dimension $\binom{n+d-1}{d}$. This space can be identified with the space $\mathbb{R}[X_{1},\ldots,X_{n}]_{d}$ of homogeneous polynomials of degree $d$ in $n$ variables, where any $T\in\text{Sym}^{d}(\mathbb{R}^{n})$ corresponds uniquely to the polynomial $\langle T,X^{\otimes d}\rangle_{F}$ with $X=(X_{1},\ldots,X_{n})^{T}$. Finally, recall that a symmetric nonzero tensor $T\in{\rm Sym}^{d}(\mathbb{R}^{n})$ has rank $1$ if $T=\lambda v^{\otimes d}$ for some $\lambda\in\mathbb{R}$ and $v\in\mathbb{R}^{n}$. More generally, a tensor $T$ has (real symmetric) rank $r$ if it can be written as a linear combination of $r$ rank-$1$ tensors $T=\lambda_{1}v_{1}^{\otimes d}+\ldots+\lambda_{r}v_{r}^{\otimes d}$, but not as a combination of $r-1$ rank-$1$ tensors. For $d=2$, this definition agrees with the usual notion of rank for symmetric matrices.
@@ 2 Shallow Polynomial Networks
where $\langle T,S\rangle_{F}$ is the Frobenius inner product of two tensors $T,S\in(\mathbb{R}^{n})^{\otimes d}$. We will sometimes write the weights compactly as $\mathcal{W}=(\alpha,W)$ where $\alpha\in\mathbb{R}^{r}$ and $W=\begin{bmatrix}w_{1}&\ldots&w_{r}\end{bmatrix}^{T}\in\mathbb{R}^{r\times n}$.
2.1 The functional space
We indicate the functional space of all polynomial networks of width $r$ with $\mathcal{F}_{r}$. It is clear that any network function $f_{\mathcal{W}}$ as in (3) is a polynomial of degree $d$ in the input $x\in\mathbb{R}^{n}$, so we may view $\mathcal{F}_{r}$ as a subset of the space of homogeneous polynomials $\mathbb{R}[X_{1},\ldots,X_{n}]_{d}$ or, equivalently, of the space of symmetric tensors ${\rm Sym}^{d}(\mathbb{R}^{n})$. Furthermore, the tensor $\sum_{i=1}^{r}\alpha_{i}w_{i}^{\otimes d}$ associated with $f_{\mathcal{W}}$ in (3) has rank at most $r$ and in fact every tensor of rank at most $r$ can be viewed as a shallow network of width $r$. The functional space $\mathcal{F}_{r}$ thus coincides with the set of symmetric tensors with bounded rank:
@@ 2.1 The functional space
From this identification we see that the functional spaces $\mathcal{F}_{r}$ are always semi-algebraic sets. It is also known that these sets not closed for $r>2$ [14] and we will return to this point later in this section. It is sometimes helpful to consider the Zariski closures $\overline{\mathcal{F}_{r}}$ of $\mathcal{F}_{r}$, as these are real algebraic varieties in ${\rm Sym}^{d}(\mathbb{R}^{n})$.
@@
Similar to [30], we say that $\mathcal{F}_{r}$ is thick if it has positive Lebesgue measure in ${\rm Sym}^{d}(\mathbb{R}^{n})$ (case $2$ and $3$ above); we say that that $\mathcal{F}_{r}$ is filling if $\mathcal{F}_{r}={\rm Sym}^{d}(\mathbb{R}^{n})$ (case $3$ above). We also let
$r_{\rm thick}(d,n)$ ::=:= $\min\{r\colon\mathcal{F}_{r}\mbox{ is thick}\},$ (4)
$r_{\rm fill}(d,n)$ ::=:= $\min\{r\colon\mathcal{F}_{r}\mbox{ is filling}\}.$
These two widths correspond to sharp qualitative changes in the function space. The value of $r_{\textup{thick}}(d,n)$ is known in all cases, but only an upper bound for $r_{\rm fill}(d,n)$ is known in general [8].
@@
[Alexander-Hirschovitz] If $d=2$, we have $r_{\textup{thick}}(2,n)=r_{\textup{fill}}(2,n)=d$. If $d\geq 3$, then
@@
These facts were already briefly discussed in [30], which considered a broader class of deep polynomial networks and mainly studied the transition from low-dimensional to thick functional spaces (referred to as “filling functional varieties”). Here, we restrict attention to shallow networks but study the nonconvex optimization landscape of these models, a topic not addressed in [30].
@@
If $n=2$ (two-dimensional input), then $\mathcal{F}_{r}$ is a family of homogeneous polynomials in $\mathbb{R}[X_{1},X_{2}]_{d}\cong\mathbb{R}^{d+1}$, or equivalently, a family of non-homogeneous polynomials in $\mathbb{R}[Z]_{\leq d}$. Assuming $d\geq 3$, then Theorem 1 states that $\mathcal{F}_{r}$ is a full-dimensional subset of $\mathbb{R}^{d+1}$ if and only if $r\geq r_{\rm thick}(d,2)=\lceil\frac{d+1}{2}\rceil$. In this case, it is also known that $r_{\rm fill}(d,2)=d$ [15, Proposition 2.1]. In the simplest case of $d=3,n=2$, then we have that $\mathcal{F}_{3}$ contains all cubic polynomials in one variable, while $\mathcal{F}_{2}$ can be characterized explicitly as the set of cubic polynomials that have exactly one real root.
@@
The notions of filling and thick function spaces are also applicable to networks with non-polynomial activations (e.g., ReLU). Indeed, while the ambient functional spaces for these networks are infinite-dimensional, a finite-dimensional embedding space always exists in the practical setting of empirical risk minimization (ERM), since functions are mapped to predictions on a finite set of training data.
@@
We first briefly describe the fibers of $\tau_{r}$ (the preimages of single points). This can also be viewed as characterizing the symmetries of the parameters in the model. We note that the parameters of any shallow network as in (3) has a set of “trivial” symmetries, namely scalings among layers and permutations among neurons. More precisely, if $D\in\mathbb{R}^{r\times r}$ is an invertible diagonal matrix and $P\in\mathbb{Z}^{r\times r}$ is a permutation matrix, then any pair of parameters $\mathcal{W}=(\alpha,W)$ and $\mathcal{W}^{\prime}=(\alpha^{\prime},W^{\prime})$ that are related by $\alpha^{\prime}=PD^{-d}\alpha$ and $W^{\prime}=PDW$ satisfy $\tau_{r}(\mathcal{W})=\tau_{r}(\mathcal{W}^{\prime})$ (i.e., $f_{\mathcal{W}}=f_{\mathcal{W}^{\prime}}$). Classical results on the identifiability of tensor decomposition imply that in many cases there are no other symmetries.
@@
Assume that $d\geq 3$, $r<r_{\rm thick}(r,d)$ and that $(d,n,r)$ is none of $(6,3,9)$, $(4,4,8)$, $(3,6,9)$. If $\mathcal{W}\in\mathbb{R}^{r}\times\left(\mathbb{R}^{n}\right)^{r}$ is Zariski-generic (i.e., belongs to an an appropriate open dense set of the parameter space) and $\mathcal{W}^{\prime}\in\mathbb{R}^{r}\times\left(\mathbb{R}^{n}\right)^{r}$ is arbitrary, then $\tau_{r}(\mathcal{W})=\tau_{r}(\mathcal{W}^{\prime})$ holds if and only if $\mathcal{W}$ and $\mathcal{W}^{\prime}$ are related by trivial symmetries.
@@
This follows from [12, Theorem 1] and [11, Remark 4]. ∎
We note that polynomial networks with quadratic activations ($d=2$) behave differently, since the parameters of the model have always more symmetries in addition to the trivial ones (for any $r$). To see this, consider $(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ such that $\alpha=(1,\ldots,1,-1,\ldots,-1)$, where $p$ elements $1$’s and $q$ elements are $-1$’s (any full rank matrix in ${\rm Sym}^{2}(\mathbb{R}^{n})$ has an element of this form in its fiber). Then we have that $W^{T}diag(\alpha)W=\tau_{r}(\alpha,W)=\tau_{r}(\alpha,MW)$ for any $M\in\mathbb{R}^{r\times r}$ that belongs to the indefinite orthogonal group $O(p,q)=\{M\colon M^{T}diag(\alpha)M=diag(\alpha)\}$. This set is a manifold (Lie group) of dimension $r(r-1)/2$ that contains $O(p)\times O(q)$. In the Appendix, we give a description of the topologically connected components of the fiber for networks with quadratic activations (Theorem 37).
We also remark that, in general, optimization landscapes in parameter space and function space are related but distinct [35]. Our paper derives results in both spaces.

### v2

@@ Abstract
We study shallow neural networks with monomial activations and output dimension one. The function space for these models can be identified with a set of symmetric tensors with bounded rank. We describe general features of these networks, focusing on the relationship between width and optimization. We then consider teacher-student problems, which can be viewed as problems of low-rank tensor approximation with respect to non-standard inner products that are induced by the data distribution. In this setting, we introduce a teacher-metric data discriminant which encodes the qualitative behavior of the optimization as a function of the training data distribution. Finally, we focus on networks with quadratic activations, presenting an in-depth analysis of the optimization landscape. In particular, we present a variation of the Eckart-Young Theorem characterizing all critical points and their Hessian signatures for teacher-student problems with quadratic networks and Gaussian training data.
@@ 1 Introduction
One of the main challenges in the theory of deep learning is to explain why neural networks can be successfully trained in practice despite the fact that their loss is highly non-convex. In recent years, researchers have argued that this behavior is mainly a consequence of overparameterization, that is, the fact that models have a large number of parameters compared to the number of training samples. However, while the benefit of overparameterization has been demonstrated in many settings, precise theoretical results often rely on unrealistic assumptions, such as infinite neurons [chizat2018, NEURIPS2018_5a4be1fa, pmlr-v99-mei19a], or only apply to very restricted models, such as linear networks [kawaguchi2016]. Additionally, several works provide sufficient conditions for the number of parameters needed for favorable optimization [du2018d, venturi2018spurious], but they do not consider important information such as the training data distribution or the initialization of gradient descent.
@@
In this paper, we seek a more precise understanding of the conditions under which local minima pose a challenge to gradient descent. Our general setting is based on shallow polynomial networks, that are functions $f_{\mathcal{W}}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ of the form
$f_{\mathcal{W}}(x)=\sum_{i=1}^{r}\alpha_{i}(w_{i}\cdot x)^{d},\qquad\mathcal{W}=(\alpha,w_{1},\ldots,w_{{r}})\in\mathbb{R}^{r}\times(\mathbb{R}^{n})^{{r}},$ (1)
@@
where $d\in\mathbb{N}$. We believe that shallow polynomial networks capture many essential features of the optimization properties for more general networks — e.g., with ReLU activations — while also being very mathematically appealing. In particular, shallow polynomial networks are intimately related to symmetric tensor decompositions: the set $\mathcal{F}_{r}$ of all shallow networks of width at most $r$ can be identified with the set of symmetric tensors of rank at most $r$. Though this connection has been noted in prior works [venturi2018spurious, kileel2019expressive], its full implications for optimization have not yet been studied in detail. Our focus on algebraic networks aligns with recent work investigating neural network architectures using algebraic-geometric tools [kohn2024function, kohn2022geometry, henry2024geometry, Kubjas_2024, finkel2024activation]. It is worth noting that the fact that polynomial activations are not universal approximators does not represent a limitation, as our study is centered on non-convex optimization. Moreover, polynomial networks of varying degrees are, in fact, universal approximators, as a consequence of the Stone-Weierstrass theorem (see also [yu2021arbitrary])
@@
We begin by revisiting the connection between tensors and shallow networks, describing properties of tensors that offer insight into these models. A key aspect that emerges from this relationship is the existence of three distinct regimes for the network width $r$: 1) for small $r$, the set $\mathcal{F}_{r}$ of shallow networks is a low-dimensional subset of its ambient space; 2) for intermediate $r$, the set $\mathcal{F}_{r}$ is a full-dimensional (semi-algebraic) set and yet strictly contained in its ambient space; 3) for large $r$, the set $\mathcal{F}_{r}$ “fills” the space of all symmetric tensors completely. These settings lead to very different behaviors when optimizing training losses, with the filling regime being the “easiest” in a precise sense. The Alexander-Hirschovitz Theorem [alexander1995] from the theory of tensors effectively classifies the sharp transitions between these regimes for all input dimensions $n$ and activation degrees $d$.
@@
After discussing general properties of polynomial networks, we focus on teacher-student training problems [loureiro2021learning, saglietti2022analytical]. For any $r,s\geq 1$, we consider a loss function of the form
@@
where we assume that the “teacher” model $\mathcal{V}$ is fixed and that we optimize “student” parameters $\mathcal{W}$. Here $\|\cdot\|$ denotes a functional norm induced by an inner product, such as an $L^{2}$ inner product $\langle f,g\rangle=\mathbb{E}_{x\sim\mathcal{D}}[f(x)g(x)]$ for some distribution of inputs $\mathcal{D}$. Minimizing (2) is equivalent to finding the optimal low-rank approximation of a given symmetric tensor, where distance is measured according to the chosen inner product in tensor space. For $L^{2}$ inner products, this distance may be expressed in terms of a finite set of moments of the input data distribution, and this applies to both discrete and continuous distributions. We explain this viewpoint in detail and describe examples of inner products arising from different classes of distributions. We also connect our setting to the broader problem of computing the distance function of points to manifolds or varieties. For that problem, it is known that qualitative changes in the optimization arise when the target point crosses a certain locus – known as ED discriminant [draisma2013] or the focal locus [milnor2016morse] – associated with the algebraic variety. We adapt this notion to define a teacher-metric data discriminant, which accounts for variations not only in the teacher but also in the metric used for computing distances. Intuitively, this discriminant encodes the qualitative behavior of the optimization landscape as a function of the training data distribution. This is a novel tool for understanding non-convex optimization problems in machine learning, which we hope to investigate further in a separate publication.
@@
When the weights of the teacher model have special symmetries (invariance to certain permutations), critical points of (2) have been identified as examples of symmetry breaking [arjevani2021symmetry] (see generally [arjevani2024symmetry]) under the class of inner products considered in this work, enabling the construction of infinite families of critical points and the characterization of their Hessian spectrum [arjevani2023symmetry]. On the other hand, when symmetry is absent, the optimization landscape for low-rank tensor approximation is not fully understood. Existing results only consider the Frobenius norm and are typically restricted to rank-$1$ approximation [ge2017optimization, draisma2018, anandkumar2014, kileel2021landscape]. Even in the simpler setting of low-rank matrix approximation, the behavior under weighted norms also remains challenging [gillis2011low]. For this reason, after some general results for arbitrary activation degrees, we conduct a detailed study for the particular case of quadratic activations. Shallow networks with quadratic activations have been the subject of several papers [du2018d, venturi2018spurious, soltanolkotabi2018, gamarnik2020, mannelli2020]. Compared to these works, our analysis provides a more general and detailed picture of the loss landscape. In particular, we provide a characterization of all critical points and their signatures in the case of Gaussian (or rotationally invariant) data distributions. Perhaps surprisingly, we also find that the teacher-student loss for “generic” data distributions behaves very differently, and leads to many more critical points than the Gaussian case. This result illustrates potential dangers of theoretical works focusing exclusively on Gaussian data. Throughout the paper, tools of applied algebraic geometry allow us to achieve novel analysis for shallow networks.
Summary of main results.
Our main contributions can be summarized as follows:
  • Function space of shallow polynomial networks: We characterize the function space $\mathcal{F}_{r}$ of shallow polynomial networks through its connection with spaces of low-rank tensors. In particular, we identify three distinct geometric regimes—low-dimensional, thick, and filling—as a function of width $r$ (Section 2.1). These regimes correspond to sharply different optimization behaviors, and their boundaries are determined by the Alexander-Hirschowitz Theorem.
  • Geometry of loss landscapes: We analyze the structure of the loss landscape, identifying different types of unfavorable behavior—namely, “bad minima” and “spurious valleys”—which may exist even in very wide networks. We also prove the existence of local minima with basins of positive Lebesgue measure (Example 15, Corollary 17), refining and extending prior work [venturi2018spurious].
  • Teacher-student problems as low-rank tensor approximation: We formulate teacher-student training problems as low-rank tensor approximation under nonstandard inner products induced by the data distribution (Section 3.1). These inner products are determined by the moments of the distribution, and we provide examples arising from various distribution families.
  • Data discriminant: We introduce the teacher-metric data discriminant, a generalization of the focal locus (or ED discriminant), which captures qualitative changes in the optimization landscape due to variations in both the teacher model and the data distribution (Theorem 26, Theorem 27). We present explicit examples with symbolic computations using Macaulay2 (Section 3.3).
  • Landscape characterization for quadratic activations: We prove versions of the Eckart-Young Theorem for teacher-student problems with quadratic activations and Frobenius (Theorem 35) or Gaussian (Theorem 36) norms. We prove that for other norms, including those arising from general i.i.d. distributions, there are generally exponentially more critical points (Theorem 37).
@@
The Frobenius inner product of two tensors $T$ and $S$ of the same shape is given by $\langle T,S\rangle_{F}=\sum_{(i_{1},\ldots,i_{d})}T_{i_{1}\ldots i_{d}}S_{i_{1}\ldots i_{d}}$. If $T=v_{1}\otimes\cdots\otimes v_{d}$ and $S=w_{1}\otimes\cdots\otimes w_{d}$, then $\langle T,S\rangle_{F}=(v_{1}\cdot w_{1})\cdots(v_{d}\cdot w_{d})$, and in particular, $\langle v^{\otimes d},w^{\otimes d}\rangle_{F}=(v\cdot w)^{d}$. A tensor $T\in\mathbb{R}^{n}\otimes\cdots\otimes\mathbb{R}^{n}$ (with $d$ factors) is symmetric if it is invariant under any permutation of indices, i.e., $T_{i_{1},\ldots,i_{d}}=T_{i_{\sigma(1)},\ldots,i_{\sigma(d)}}$ for any $\sigma\in\mathfrak{S}_{d}$, the symmetric group of degree $d$. We denote the space of symmetric tensors of order $d$ on $\mathbb{R}^{n}$ by $\text{Sym}^{d}(\mathbb{R}^{n})$, a vector space of dimension $\binom{n+d-1}{d}$. This space can be identified with the space $\mathbb{R}[X_{1},\ldots,X_{n}]_{d}$ of homogeneous polynomials of degree $d$ in $n$ variables, where any $T\in\text{Sym}^{d}(\mathbb{R}^{n})$ corresponds uniquely to the polynomial $\langle T,X^{\otimes d}\rangle_{F}$ with $X=(X_{1},\ldots,X_{n})^{T}$. Finally, recall that a symmetric nonzero tensor $T\in{\rm Sym}^{d}(\mathbb{R}^{n})$ has rank $1$ if $T=\lambda v^{\otimes d}$ for some $\lambda\in\mathbb{R}$ and $v\in\mathbb{R}^{n}$. More generally, a tensor $T$ has (real symmetric) rank $r$ if it can be written as a linear combination of $r$ rank-11 tensors $T=\lambda_{1}v_{1}^{\otimes d}+\ldots+\lambda_{r}v_{r}^{\otimes d}$, but not as a combination of r1r-1 rank-11 tensors. For $d=2$, this definition agrees with the usual notion of rank for symmetric matrices.
@@ 2 Shallow Polynomial Networks
Section overview. We introduce shallow polynomial networks and their connection to symmetric tensor decompositions. We classify expressivity regimes—low-dimensional, thick, and filling—based on network width, and analyze how these transitions influence the loss landscape. In particular, we describe how critical points, bad minima, and spurious valleys depend on width, extending prior work [venturi2018spurious] and showing that unfavorable critical points can occur even in wide networks, sometimes with basins of attraction of positive measure.
@@
where $\langle T,S\rangle_{F}$ is the Frobenius inner product of two tensors $T,S\in(\mathbb{R}^{n})^{\otimes d}$. Throughout the paper, we assume that the network has scalar output, i.e., output dimension one. We will sometimes write the weights compactly as $\mathcal{W}=(\alpha,W)$ where $\alpha\in\mathbb{R}^{r}$ and $W=\begin{bmatrix}w_{1}&\ldots&w_{r}\end{bmatrix}^{T}\in\mathbb{R}^{r\times n}$.
2.1 The function space
We indicate the function space of all polynomial networks of width $r$ with $\mathcal{F}_{r}$. It is clear that any network function $f_{\mathcal{W}}$ as in (3) is a polynomial of degree $d$ in the input $x\in\mathbb{R}^{n}$, so we may view $\mathcal{F}_{r}$ as a subset of the space of homogeneous polynomials $\mathbb{R}[X_{1},\ldots,X_{n}]_{d}$ or, equivalently, of the space of symmetric tensors ${\rm Sym}^{d}(\mathbb{R}^{n})$. Furthermore, the tensor $\sum_{i=1}^{r}\alpha_{i}w_{i}^{\otimes d}$ associated with $f_{\mathcal{W}}$ in (3) has rank at most $r$ and in fact every tensor of rank at most $r$ can be viewed as a shallow network of width $r$. The function space $\mathcal{F}_{r}$ thus coincides with the set of symmetric tensors with bounded rank:
@@ 2.1 The function space
From this identification we see that the function spaces $\mathcal{F}_{r}$ are always semi-algebraic sets. It is also known that these sets are not closed for $r>2$ [comon2008] and we will return to this point later in this section. It is sometimes helpful to consider the Zariski closures $\overline{\mathcal{F}_{r}}$ of $\mathcal{F}_{r}$, as these are real algebraic varieties in ${\rm Sym}^{d}(\mathbb{R}^{n})$.
@@
Similar to [kileel2019expressive], we say that $\mathcal{F}_{r}$ is thick if it has positive Lebesgue measure in ${\rm Sym}^{d}(\mathbb{R}^{n})$ (case $2$ and $3$ above); we say that $\mathcal{F}_{r}$ is filling if $\mathcal{F}_{r}={\rm Sym}^{d}(\mathbb{R}^{n})$ (case $3$ above). We also let
$r_{\rm thick}(d,n)$ = $\min\{r\colon\mathcal{F}_{r}\mbox{ is thick}\},$ (4)
$r_{\rm fill}(d,n)$ = $\min\{r\colon\mathcal{F}_{r}\mbox{ is filling}\}.$
These two widths correspond to sharp qualitative changes in the function space. The value of $r_{\textup{thick}}(d,n)$ is known in all cases, but only an upper bound for $r_{\rm fill}(d,n)$ is known in general [bernardi2018real].
@@
If $d=2$, we have $r_{\textup{thick}}(2,n)=r_{\textup{fill}}(2,n)=n$. If $d\geq 3$, then
@@
These facts were already briefly discussed in [kileel2019expressive], which considered a broader class of deep polynomial networks and mainly studied the transition from low-dimensional to thick function spaces (referred to as “filling functional varieties”). Here, we restrict attention to shallow networks but study the nonconvex optimization landscape of these models, a topic not addressed in [kileel2019expressive].
@@
If $n=2$ (two-dimensional input), then $\mathcal{F}_{r}$ is a family of homogeneous polynomials in $\mathbb{R}[X_{1},X_{2}]_{d}\cong\mathbb{R}^{d+1}$, or equivalently, a family of non-homogeneous polynomials in in a single variable. Assuming $d\geq 3$, then Theorem 1 states that $\mathcal{F}_{r}$ is a full-dimensional subset of $\mathbb{R}^{d+1}$ if and only if $r\geq r_{\rm thick}(d,2)=\lceil\frac{d+1}{2}\rceil$. In this case, it is also known that $r_{\rm fill}(d,2)=d$ [comon2012typical, Proposition 2.1]. In the simplest case of $d=3,n=2$, then we have that $\mathcal{F}_{3}$ contains all cubic polynomials in one variable, while $\mathcal{F}_{2}$ can be characterized explicitly as the set of cubic polynomials that have exactly one real root.
@@
The notions of filling and thick function spaces are also applicable to networks with non-polynomial activations (e.g., ReLU). Indeed, while the ambient function spaces for these networks are infinite-dimensional, a finite-dimensional embedding space always exists in the practical setting of empirical risk minimization (ERM), since functions are mapped to predictions on a finite set of training data.
@@
We first briefly describe the fibers of $\tau_{r}$ (the preimages of single points). This can also be viewed as characterizing the symmetries of the parameters in the model. We note that the parameters of any shallow network as in (3) has a set of “trivial” symmetries, namely scalings among layers and permutations among neurons. More precisely, if $D\in\mathbb{R}^{r\times r}$ is an invertible diagonal matrix and $P\in\mathbb{Z}^{r\times r}$ is a permutation matrix, then any pair of parameters $\mathcal{W}=(\alpha,W)$ and $\mathcal{W}^{\prime}=(\alpha^{\prime},W^{\prime})$ that are related by $\alpha^{\prime}=PD^{-d}\alpha$ and $W^{\prime}=PDW$ satisfies $\tau_{r}(\mathcal{W})=\tau_{r}(\mathcal{W}^{\prime})$ (i.e., $f_{\mathcal{W}}=f_{\mathcal{W}^{\prime}}$). Classical results on the identifiability of tensor decomposition imply that in many cases there are no other symmetries.
@@
Assume that $d\geq 3$, $r<r_{\rm thick}({d},{n})$ and that $(d,n,r)$ is none of $(6,3,9)$, $(4,4,8)$, $(3,6,9)$. If $\mathcal{W}\in\mathbb{R}^{r}\times\left(\mathbb{R}^{n}\right)^{r}$ is Zariski-generic (i.e., belongs to an appropriate open dense set of the parameter space) and $\mathcal{W}^{\prime}\in\mathbb{R}^{r}\times\left(\mathbb{R}^{n}\right)^{r}$ is arbitrary, then $\tau_{r}(\mathcal{W})=\tau_{r}(\mathcal{W}^{\prime})$ holds if and only if $\mathcal{W}$ and $\mathcal{W}^{\prime}$ are related by trivial symmetries.
@@
This follows from [chiantini2017generic, Theorem 1] and [chiantini2017effective, Remark 4]. ∎
We note that polynomial networks with quadratic activations ($d=2$) behave differently, since the parameters of the model have always more symmetries in addition to the trivial ones (for any $r$). To see this, consider $(\alpha,W)\in\mathbb{R}^{r}\times\mathbb{R}^{r\times n}$ such that $\alpha=(1,\ldots,1,-1,\ldots,-1)$, where $p$ elements are $1$’s and $q$ elements are 1-1’s (any full rank matrix in ${\rm Sym}^{2}(\mathbb{R}^{n})$ has an element of this form in its fiber). Then we have that $W^{T}\operatorname{diag}(\alpha)W=\tau_{r}(\alpha,W)=\tau_{r}(\alpha,MW)$ for any $M\in\mathbb{R}^{r\times r}$ that belongs to the indefinite orthogonal group $O(p,q)=\{M\colon M^{T}\operatorname{diag}(\alpha)M=\operatorname{diag}(\alpha)\}$. This set is a manifold (Lie group) of dimension r(r1)/2r(r-1)/2 that contains $O(p)\times O(q)$. In the Appendix, we give a description of the topologically connected components of the fiber for networks with quadratic activations (Theorem 38).
We also remark that, in general, optimization landscapes in parameter space and function space are related but distinct [trager2019pure, levin2024effect]. Our paper derives results in both spaces.